# VIMD: Monocular Visual-Inertial Motion and Depth Estimation

### v2

@@ Abstract
Accurate and efficient dense metric depth estimation is crucial for 3D visual perception in robotics and XR. In this paper, we develop a monocular visual-inertial motion and depth (VIMD) learning framework to estimate dense metric depth by leveraging accurate and efficient MSCKF-based monocular visual-inertial motion tracking. At the core the proposed VIMD is to exploit multi-view information to iteratively refine per-pixel scale, instead of globally fitting an invariant affine model as in the prior work. The VIMD framework is highly modular, making it compatible with a variety of existing depth estimation backbones. We conduct extensive evaluations on the TartanAir and VOID datasets and demonstrate its zero-shot generalization capabilities on the AR Table dataset. Our results show that VIMD achieves exceptional accuracy and robustness, even with extremely sparse points—as few as 10-20 metric depth points per image. This makes the proposed VIMD a practical solution for deployment in resource-constrained settings, while its robust performance and strong generalization capabilities offer significant potential across a wide range of scenarios.
@@ I Introduction
Figure 1: A motivation result: Invariant affinity is not sufficient to scale monocular depth prediction. Per-frame least-squares fitting of a global scale and offset (every 5th frame) using VOID data [2]. Clearly, the offset varies more than the scale, indicating that a single global affine model cannot reliably align predictions over time.
Although recent advances in monocular depth prediction have achieved high-quality relative depth estimation [3, 4], these methods still lack the ability to produce outputs with absolute metric scale. A common approach assumes an invariant affinity model and aligns the predicted depth to the sparse VIO depth via per-frame least-squares fitting for a global scale and offset (e.g., see [5]). However, as shown in Fig. 1, the fitted scale and offset vary significantly across frames—particularly the offset, which exhibits much larger fluctuations than the scale. This temporal instability indicates that the predicted relative depths are not truly affine with respect to the metric depths; that is, they cannot be accurately modeled by a single global scale and offset per frame without unmodeled errors. Besides the unreliable per-frame global fitting, the optimization may become poorly conditioned when sparse points are unevenly distributed or noisy, leading to sensitivity in the fitted parameters and potential error amplification in low-density regions. It has been empirically observed that predicting offset along with scale can hurt depth accuracy [6].
From Fig. 1, it is clear that the offset exhibits much higher variance across frames than the scale, implying that offsets may also vary more significantly across pixels within a single frame (due to factors like viewpoint changes, occlusions, or non-uniform scene structure), thus making them harder to learn reliably at the per-pixel level compared to the more stable scale. Motivated by this observation, we propose to exploit multi-view constraints and learn to optimize iteratively a per-pixel scale, formulated as learning approach that moves beyond the global fitting affine parameters. In particular, the proposed monocular Visual-Inertial Motion and Depth (VIMD) leverages an accurate and efficient MSCKF-based visual-inertial motion estimation module to form multi-view geometric constraints and then refines and predicts dense depth and its uncertainty. As a result, the proposed VIMD improves accuracy and robustness in low-density depth regions. Fig. 2 shows some exemplar performance of the proposed VIMD in both outdoor and indoor scenes, producing high-quality dense metric depths.
@@
Figure 2: Overall performance of the proposed VIMD: High-quality dense metric depths in both outdoor and indoor scenes.
@@ IV-A Evaluation on VOID
More interestingly, beyond absolute accuracy, Tab. II shows the comparative results under different levels of sparse point reduction. Even when the number of points is reduced by 80%, our method consistently outperforms GA and SML across all metrics. Notably, the relative advantage over SML increases as the point density decreases, highlighting the robustness of our approach. This robustness arises from two key factors. First, our method leverages multi-view information through pose integration, which provides additional geometric constraints that compensate for the reduced spatial coverage of sparse points. Second, the ConvGRU module learns to iteratively refine the scale, enabling the network to progressively optimize depth estimates across frames. By combining multi-view cues with recurrent scale refinement, our model is able to maintain both accuracy and stability, even during low density sparse points.
@@ V Conclusions and Future Work
In this paper, we have developed a novel monocular visual-inertial motion and depth (VIMD) model to learn dense metric depth from the minimal sensing platform of a monocular camera and IMU. At the core of our approach is to exploit multi-view information within a learning-based framework to iteratively optimize per-pixel scale. Additionally, the prediction of uncertainty yields more reliable depth for downstream applications and supervision at full image resolution enhances scale accuracy. As a result, the proposed VIMD model is robust in scenarios with sparse features and demonstrates generalization ability, including zero-shot performance on real-world datasets.
Nevertheless, we did observe that the cost-geometry optimization remains sensitive to textureless regions and pose drift, which can limit performance in challenging environments. Therefore, for future work, we will extend this model into a full-fledged efficient SLAM by training in an end-to-end fashion that jointly optimizes both pose and scale. Note that the proposed VIMD model can be used as a prior to Gaussian-Splatting (GS)-SLAM [29] for monocular language photorealistic mapping.

### v3

@@ Abstract
Accurate and efficient dense metric depth estimation is crucial for 3D visual perception in robotics and XR. In this paper, we develop a monocular visual-inertial motion and depth (VIMD) learning framework to estimate dense metric depth by leveraging accurate and efficient MSCKF-based monocular visual-inertial 3D motion tracking. At the core the proposed VIMD is to exploit multi-view information to iteratively refine per-pixel scale, instead of globally fitting an invariant affine model as in the prior work. The VIMD framework is highly modular, making it compatible with a variety of existing depth estimation backbones. We conduct extensive evaluations on the TartanAir and VOID datasets and demonstrate its zero-shot generalization capabilities on the AR Table dataset. Our results show that VIMD achieves exceptional accuracy and robustness, even with extremely sparse points—as few as 10-20 metric depth points per image. This makes the proposed VIMD a practical solution for deployment in resource-constrained settings, while its robust performance and strong generalization capabilities offer significant potential across a wide range of scenarios.
@@ I Introduction
Figure 1: A motivation result: A single global affine model fails for metric depth recovery from monocular predictions. Left: depth error after per-frame global scale alignment. Right: per-pixel ideal scale from ground truth, revealing strong spatial variation.
Although recent advances in monocular depth prediction achieve high-quality relative depth estimation [19, 2], these methods still lack the ability to produce depth estimates with absolute metric scale. A common approach assumes an invariant affine relationship and aligns the predicted depth to sparse visual–inertial odometry (VIO) depth measurements via per-frame least-squares fitting of a global scale and offset (e.g., see [14, 15]). While simple and computationally efficient, this approach implicitly assumes that a single global affine transformation is sufficient to correct the predicted depth across the entire image.
As shown in Fig. 1, globally aligned depth (left) exhibits noticeable errors that increase as fewer sparse depth points are available (see Tab. II). To further analyze this, we compute the ideal scale at each pixel using ground-truth depth and the globally aligned depth. The resulting scale map (right) shows strong spatial variation, indicating that predicted and metric depth are not globally affine. As a result, enforcing a single global scale (or scale and offset) introduces systematic, spatially varying errors that per-frame fitting cannot correct.
Beyond the unreliability of per-frame global fitting, the optimization becomes poorly conditioned when sparse depth points are low. This leads to increased sensitivity of the fitted parameters and potential error amplification, particularly in low-density regions. Furthermore, it has been empirically observed that jointly predicting an offset along with scale can degrade depth accuracy [24].
@@
Figure 2: Illustration of the proposed VIMD learning high-quality dense metric depths in both outdoor and indoor scenes.
Motivated by these observations, we investigate how to better leverage sparse VIO depth measurements and camera motion to obtain efficient, robust, and accurate metric depth for 3D spatial perception. Our contributions are:
  • We propose VIMD, a modular monocular visual–inertial framework that leverages multi-view geometric constraints from an efficient MSCKF-based visual–inertial motion estimation to learn to iteratively refine per-pixel scale for dense metric depth estimation, moving beyond per-frame global affine alignment.
  • We develop a lightweight GRU-based iterative depth refinement model, in which reference frames are warped to the target frame using the predicted depth and VIO poses, and the scale is iteratively refined using a ConvGRU to predict the final depth and uncertainty.
  • We demonstrate that VIMD improves depth accuracy and robustness in low-density depth regions (see Tab. II) and produces high-quality dense metric depth maps in both indoor and outdoor scenes (see Fig. 2). Extensive experiments further show strong zero-shot generalization and real-time performance.
The rest of the paper is structured as follows: After reviewing the literature in the next section, we present in detail the proposed metric dense depth learning only from a monocular camera and IMU in Section III. The extensive experimental results are presented in Section IV. We conclude the paper in Section V along with future research directions.
@@ IV-A Evaluation on VOID
More interestingly, beyond absolute accuracy, Tab. LABEL:tab:combined_metrics_selected shows the comparative results under different levels of sparse point reduction. Even when the number of points is reduced by 80%, our method consistently outperforms GA and SML across all metrics. Notably, the relative advantage over SML increases as the point density decreases, highlighting the robustness of our approach. This robustness arises from two key factors. First, our method leverages multi-view information through pose integration, which provides additional geometric constraints that compensate for the reduced spatial coverage of sparse points. Second, the ConvGRU module learns to iteratively refine the scale, enabling the network to progressively optimize depth estimates across frames. By combining multi-view cues with recurrent scale refinement, our model is able to maintain both accuracy and stability, even during low density sparse points.
@@ V Conclusions and Future Work
In this paper, we presented VIMD, a monocular visual–inertial motion and depth framework that recovers dense metric depth using only a monocular camera and an IMU. At the core of our approach is to leverage multi-view information within a learning-based framework to iteratively optimize per-pixel scale. Additionally, uncertainty prediction improves the reliability of the estimated depth for downstream tasks, while full-resolution supervision further enhances final depth accuracy. As a result, the VIMD is robust under fewer sparse features and generalizes well across environments, including zero-shot performance on real-world datasets.
We also identify current limitations: iterative refinement can degrade in textureless regions and under inaccurate pose estimates. As a practical mitigation, the proposed refinement module can be applied after sparse bundle adjustment, where more reliable camera poses are available. In the future work, we will extend this model into a tightly-coupled SLAM by training in an end-to-end fashion. Finally, VIMD can serve as a strong metric depth prior for downstream mapping systems, such as Gaussian-Splatting (GS)-SLAM [12] for monocular language photorealistic mapping.